The heating coil rated 1kw is used to evaporated 700g of water at 100 degree.if specific latent heat of vaporization of water is 2.e×10^6, the time taken to evaporate all the water is?
1
Expert's answer
2019-02-04T10:01:48-0500

Let's write the formula for the amount of heat required to change the state of water at

100100 ℃

to a gas:

Q=mL,Q = mL,

here,

QQ

is the amount of heat required to change the state of water at

100100 ℃

to a gas,

m=0.7kgm = 0.7 kg

is the mass of the water and

L=2106JkgL = 2 \cdot 10^6 \frac{J}{kg}

is the specific latent heat of vaporization of water.

From the other hand, we can write the formula for the amount of heat that a coil provides at the rate of

1kW1 kW

during time

tt

:

Q=Pt,Q = Pt,

here,

tt

is the time taken to evaporate all the water.

Then, we can equate both expressions and find the time taken to evaporate all the water:

mL=Pt,mL = Pt,t=mLP=0.7kg2106Jkg1000W=1400s.t = \frac{mL}{P} = \frac{0.7 kg \cdot 2 \cdot 10^6 \frac{J}{kg}}{1000 W} = 1400 s.

Answer:

t=1400st = 1400 s

.